# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 6.2.1. First approaches
Given any semi-Finsler metric $L$ with fundamental tensor $g$ and vector field $V$ lying in its domain, one can define the osculating semi-Riemannian metric $g_{V}(x):=g(x,V(x))$. Asanov [6] carried out a variational approach for such metrics leading to an equation formally similar to Einstein’s vacuum one. However, the subtleties of the osculating metrics (see for example the interpretation of Chern anisotropic connection in §2.4.5) make unclear its implications.
Miron [128] considered Finsler metrics from a general Lagrange viewpoint. Starting at the fundamental tensor $g$, a Sasaki-type metric on $A\subset TM\setminus\mathbf{0}$, which would admit a classical Einstein equation, is constructed. Even though no further physical interpretation was given, the Lagrangian theory and its physical applicability is analyzed further in Miron et al. [129].
@@ 6.2.2. Pfeifer–Wohlfarth (PW) equation
Pfeifer and Wohlfarth [146] proposed an action functional for the gravitational sector of Finsler geometry by using the Ricci scalar
$\mathscr{S}[L]:=\int_{\mathbb{P}^{+}A}g^{\alpha\beta}\,\mathrm{Ric}_{\cdot\alpha\cdot\beta}\,d\Sigma^{+}.$ (27)
(recall the dot notation for vertical derivatives in (15)), where $d\Sigma^{+}$ denotes the natural volume form on the positively projectivized tangent bundle $\mathbb{P}^{+}A$. This action is closely related to the Einstein–Hilbert functional in Riemannian geometry, and it represents the simplest choice based on the canonical curvature derived from the connection coefficients $\mathring{G}^{\mu}$.
@@
where $\mathfrak{T}$ denotes the energy–momentum distribution function on $A$, and $\mathrm{Lan}_{\mu}$ is the mean Landsberg tensor (§2.4.5). PW equation makes sense in any signature. For semi-Riemannian metrics in vacuum ($\mathfrak{T}=0$), it is equivalent to impose Ricci flatness, but not Rutz equation (26) in general.
@@
which is consistent with known physical conservation laws, such as those arising in the Liouville equation for kinetic gases. What is more, the authors conjectured that the refinement of $\mathfrak{T}$ from a perfect fluid description to a kinetic gas, combined with the effects of the Landsberg tensor, could give rise to an accelerated cosmic expansion without invoking dark energy.
6.3. Exact Vacuum Solutions of PW
@@
The solutions are typically analytic, defined on the whole $TM$ but possibly containing singularities in some directions which may lie in the causal cone.
@@ 6.3.1. $(\alpha,\beta)$-metrics and pp-waves
[82, Th. 8.8.1]. Let $F$ be any $(\alpha,\beta)$-metric such that $\alpha$ solves the classical Einstein equations in vacuum and $\beta$ is parallel respect to $\alpha$. Then, the nonlinear connection of $F$ coincides with the Levi-Civita connection of $\alpha$ and $F$ becomes a (Ricci-flat) vacuum solution to PW equation.
@@
[82, Th. 8.8.2] Let $(u,v,x^{1},x^{2})$ be local coordinates. Any $(\alpha,\beta)$-metric defined by
@@
$F=\frac{\alpha^{1+m}}{\beta^{m}},\qquad m\in\mathds{R},$ (32)
which also include the Bogoslovsky metrics (23). We refer to [145, 82] for a systematic study.
@@ 6.3.2. Unicorns and Finslerian Cosmologies
Building upon Elgendi’s unicorns in Finsler metrics with singularities, Heefer et al. [83] arrive at semi-Finsler unicorns with singularities and, remarkably, some of them are solutions of PW equations. More precisely (see formulas (10.8), (10.30) —and below— and Prop. 10.2.2 in [82]):
@@
The light cones for $F$ are equal to those of Lorentz-Minkowski if the signature of $\Phi$ is either Lorentzian (+,+,)(+,+,-) or negative (,,)(-,-,-) . However, the signature of the fundamental tensor $g$ is not Lorentzian inside the cone in these cases, [82, Prop. 10.1.4]. Indeed, for Lorentzian $\Phi$, one can consider the sets $\mathcal{S}(\Phi),\mathcal{T}(\Phi)$ containing, resp., $\Phi$-spacelike and $\Phi$-timelike directions. Then, $g$ has Lorentz signature on $\mathds{R}\times\mathcal{S}(\Phi)$ but neutral (+,+,,)(+,+,-,-) on $\mathds{R}\times\mathcal{T}(\Phi)$. For negative definite $\Phi$ the signature of $g$ is also negative definite.
This may make obscure the relativistic interpretation of such unicorns. However, they present further interesting properties to be explored in both the cosmological and purely geometric setting. Indeed, in the case of negative definite $\Phi$, the unicorns correspond to a linearly expanding (or contracting) universe with a cosmological symmetry which make them reminiscent of FLRW cosmology. They are not only conformally flat as them but also flat for curvature of the non-linear connection, that is $\mathcal{R}_{ij}^{k}=\delta_{j}N_{i}^{k}-\delta_{i}N_{j}^{k}\equiv 0$ (recall (15)). This is precisely a consequence of the non-Berwaldian character of the metric and, thus, a potentially measurable Finsler effect, beyond the scope of standard General Relativity.
@@ 6.4. Finsler Einstein-Hilbert-Palatini approach
The huge freedom introduced by the Finslerian viewpoint permits to consider a good number of functionals, depending on the focused geometric level of the structure (non-linear, anisotropic or Finsler connection), see the discussion in [151]. Anyway, the so-called Palatini [^26] It is kwown that this aproach was introduced by Einstein himself, but we use it following [100] because it is widely spread (an example of Stigler’s Law of Eponymy) and distinguishes from other related formalisms. formulation is a natural extension of the Hilbert variational approach which fits especially well to the Finsler viewpoint, developed in [100] (our main reference below). Indeed, Palatini’s promote the nonlinear connection (which determines the freely falling observers) to geometric field independent of the metric (which provides numbers from observations), as they come from a priori independent phenomena. This theory has an independent variational interest in semi-Finsler Geometry and formally recovers Rutz’s (26), a desirable property obtained by PW only when the mean Landsberg tensor vanishes. What is more, it diverges significantly of PW when $\mathrm{Lan}_{\mu}\neq 0$, as the geodesics of the connection will differ then (in particular, predicting experimentaly testable differences).
@@ 6.4.1. Palatini Formalism in Semi-Riemannian Geometry
where the metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\mu}_{\nu\rho}$ are treated as independent variables, $\mathrm{ric}_{\alpha\beta}$ is the Ricci tensor of the connection, and the metric provides the contraction and volume. A pair $(g_{\mu\nu},\Gamma^{\mu}_{\nu\rho})$ is a critical point of $\mathscr{I}_{\mathrm{EHP}}$ if and only if $g_{\mu\nu}$ satisfies the Einstein field equations and the affine $\Gamma^{\mu}_{\nu\rho}$ satisfies, in terms of Christoffel symbols $(\Gamma^{g})_{ij}^{k}$ of the Levi-Cita connection of $g_{\mu\nu}$ (see, for example [21])
$\Gamma_{ij}^{k}=(\Gamma^{g})_{ij}^{k}+\mathcal{A}_{i}\otimes\delta_{j}^{k},$ (35)
where $\mathcal{A}$ is 1-form (with no restriction) and $\delta_{j}^{k}$ is Kronecker’s delta. The unique symmetric connection in (35) is Levi-Civita one, then providing a remarkable support to standard General Relativity.
@@ 6.4.2. Palatini Approach to Finslerian Einstein Equations
$\bar{\mathscr{S}}[L,N^{\mu}_{\nu}]:=\int g^{\alpha\beta}\,\mathrm{Ric}_{\cdot\alpha\cdot\beta}\,d\Sigma^{+},$ (36)
@@ 6.4.3. Some results on the Finslerian Palatini Equations
The following sampler of results in [100] is illustrative. The first one stresses an aforementioned fundamental difference with PW (see more complete one [100, Th. D]).
@@
Equation (37) resembles Einstein’s, which reduced to the vanishing of the Ricci tensor just taking traces. Palatini vacuum reduces to the vanishing of the Ricci scalar (in the spirit of Rutz’s (26)) but the proof is not trivial because now a differential equation must be solved [100, Lemma 5.13]. Surprisingly, the result is obtained only for Lorentzian signature, because the proof is then based on the Riemannian maximum principle on the indicatrix [100, Th. C] (in the case that $L$ is Riemannian, an alternative proof could be given, based on the eigenvalues of the Laplacian on a round sphere, cf. [100, Th. 5.17]).
@@
In Lorentzian signature, the solutions $(L,N)$ of the Palatini equations (37), (38) such that $N$ is defined on the whole cone $\mathcal{C}$ of $L$ satisfy $\mathrm{Ric}=0.$
@@
As in the case of semi-Riemannian Palatini (35), given a pseudo-Finsler metric $L$, the solutions of the affine equation (38) have a fibered structure on the set of symmetric solutions (now with fiber isomorphic to the space of anisotropic 0-homogeneous 1-forms A), [100, Th. A]. However, the uniqueness of the symmetric case is not trivial. Again, we can obtain uniqueness, by using a fiberwise argument of analiticy. Notice that even non-smooth semi-Riemannian metrics are fiberwise analytic, as the metric at each point is a quadric, thus analytic. The highly original argument is valid for indefinite signature, as it will use crucially analiticity when $L$ vanishes [100, Th. B], and the solutions are assumed to be well defined therein, i.e. proper.
@@ Acknowledgments
The author acknowledges warmly D. Azagra for his advise on smoothening in the proof of Th. 3.3, F. F. Villaseñor (U. Castilla la Mancha) for his comments on the manuscript, especially sections 2.4.4 to 2.4.6 and 6 and M.A. Javaloyes (U. Murcia) for his reading of the manuscript and suggestions. Research partially supported by PID2024-156031NB-I00 (MCIN/ AEI /10.13039/501100011033), as well as the framework IMAG/ María de Maeztu, CEX2020-001105-MCIN/ AEI/ 10.13039/501100011033.

### v2

@@ 6.2.1. First approaches
Given any semi-Finsler metric $L$ with fundamental tensor $g$ and any vector field $V$ lying in its domain, one can define the osculating semi-Riemannian metric $g_{V}(x):=g(x,V(x))$. Asanov [7] carried out a variational approach for such metrics leading to an equation formally similar to Einstein’s vacuum one. However, the subtleties of the osculating metrics (see for example the interpretation of Chern anisotropic connection in §2.4.5) make unclear its implications.
Miron [138] considered Finsler metrics from a general Lagrange viewpoint. Starting at the fundamental tensor $g$, a Sasaki-type metric on $A\subset TM\setminus\mathbf{0}$, which would admit a classical Einstein equation, is constructed. No further physical interpretation was given then, anyway, the Lagrangian theory and its physical applicability is analyzed further by Miron et al. [139].
@@ 6.2.2. Pfeifer–Wohlfarth (PW) equation
Pfeifer and Wohlfarth [159] proposed an action functional for the gravitational sector of Finsler geometry by integrating the Ricci scalar on the indicatrix or, equivalently [108, §3],
$\mathscr{S}[L]:=\int_{D\subset\mathbb{P}^{+}A}g^{\alpha\beta}\,\mathrm{Ric}_{\cdot\alpha\cdot\beta}\,d\Sigma^{+},$ (30)
where $d\Sigma^{+}$ denotes the natural volume form on the positively projectivized tangent bundle $\mathbb{P}^{+}A$. This action is closely related to the Einstein–Hilbert functional in Riemannian geometry, and represents the simplest choice based on the canonical curvature derived from the connection coefficients $\mathring{G}^{\mu}$.
@@
(recall the dot notation for vertical derivatives in (17)) where $\mathfrak{T}$ denotes the energy–momentum distribution function on $A$, and $\mathrm{Lan}_{\mu}$ is the mean Landsberg tensor (§2.4.5). PW equation makes sense in any signature. For semi-Riemannian metrics in vacuum ($\mathfrak{T}=0$), it is equivalent to impose Ricci flatness, but not to Rutz equation (29) in general.
@@
which is consistent with known physical conservation laws, such as those arising in the Liouville equation for kinetic gases. What is more, the authors conjectured that the refinement of $\mathfrak{T}$ from a perfect fluid description to a kinetic gas, combined with the effects of the Landsberg tensor, could give rise to an accelerated cosmic expansion without invoking dark energy [94, 158].
6.3. Exact Vacuum Solutions to PW’s
@@
The solutions are typically analytic, defined on the whole $TM$ but possibly containing singularities in some directions, which may lie in the causal cone.
@@ 6.3.1. $(\alpha,\beta)$-metrics and pp-waves
[89, Th. 8.8.1]. Let $F$ be any $(\alpha,\beta)$-metric such that $\alpha$ solves the classical Einstein equations in vacuum and $\beta$ is parallel respect to $\alpha$. Then, the nonlinear connection of $F$ coincides with the Levi-Civita connection of $\alpha$ and $F$ becomes a (Ricci-flat) vacuum solution to PW equation.
@@
[89, Th. 8.8.2] Let $(u,v,x^{1},x^{2})$ be local coordinates. Any $(\alpha,\beta)$-metric defined by
@@
$F=\frac{\alpha^{1+m}}{\beta^{m}},\qquad m\in\mathds{R}.$ (35)
These metrics also include Bogoslovsky’s in (26), and we refer to [157, 89] for a systematic study.
@@ 6.3.2. Unicorns and Finslerian Cosmologies
Building upon Elgendi’s unicorns in Finsler metrics with singularities, Heefer et al. [90] arrive at semi-Finsler unicorns with singularities and, remarkably, some of them are solutions of PW equations. More precisely (see [89], formulas (10.8), (10.30) —and below— and Prop. 10.2.2):
@@
The light cones for $F$ are equal to those of Lorentz-Minkowski if the signature of $\Phi$ is either Lorentzian (+,+,)(+,+,-) or negative (,,)(-,-,-) . However, the signature of the fundamental tensor $g$ is not Lorentzian inside the cone in these cases, [89, Prop. 10.1.4]. Indeed, for Lorentzian $\Phi$, one can consider the sets $\mathcal{S}(\Phi),\mathcal{T}(\Phi)$ containing, resp., $\Phi$-spacelike and $\Phi$-timelike directions. Then, $g$ has Lorentz signature on $\mathds{R}\times\mathcal{S}(\Phi)$ but neutral (+,+,,)(+,+,-,-) on $\mathds{R}\times\mathcal{T}(\Phi)$. For negative definite $\Phi$ the signature of $g$ is also negative definite.
These properties may obscure the relativistic interpretation of such unicorns. However, they present further interesting properties to be explored in both the cosmological and purely geometric setting. Indeed, in the case of negative definite $\Phi$, the unicorns correspond to a linearly expanding (or contracting) universe with a cosmological symmetry which make them reminiscent of FLRW cosmology. They are not only conformally flat as FLRW’s but also flat for the curvature of the non-linear connection, that is $\mathcal{R}_{ij}^{k}=\delta_{j}N_{i}^{k}-\delta_{i}N_{j}^{k}\equiv 0$ (recall (17)). This is precisely a consequence of the non-Berwaldian character of the metric and, thus, a potentially measurable Finsler effect, beyond the scope of standard General Relativity.
@@ 6.4. Finsler Einstein-Hilbert-Palatini approach
The huge freedom introduced by the Finslerian viewpoint permits to consider a good number of functionals, depending on the focused geometric level of the structure (non-linear, anisotropic or Finsler connection), see the discussion in [165]. Anyway, the so-called Palatini [^26] It is kwown that this aproach was introduced by Einstein himself, but we use it following [108] because it is widely spread (an example of Stigler’s Law of Eponymy) and distinguishes from other related formalisms. formulation is a natural extension of the Hilbert variational approach which fits especially well to the Finsler viewpoint, developed in [108] (our main reference below). Indeed, Palatini’s promote the nonlinear connection (which determines the freely falling observers) to a geometric field independent of the metric (which provides numbers from observations), as they come from a priori independent phenomena. This theory has an independent variational interest in semi-Finsler Geometry and formally recovers Rutz’s (29), a desirable property obtained by PW only when the mean Landsberg tensor vanishes. What is more, it diverges significantly of PW when $\mathrm{Lan}_{\mu}\neq 0$, as the geodesics of the connection will differ then (in particular, predicting experimentaly testable differences).
@@ 6.4.1. Palatini Formalism in Semi-Riemannian Geometry
where the metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\mu}_{\nu\rho}$ are treated as independent variables, $\mathrm{ric}_{\alpha\beta}$ is the Ricci tensor of the connection, and the metric provides the contraction and volume. A pair $(g_{\mu\nu},\Gamma^{\mu}_{\nu\rho})$ is a critical point of $\mathscr{I}_{\mathrm{EHP}}$ if and only if $g_{\mu\nu}$ satisfies the Einstein field equations and the affine $\Gamma^{\mu}_{\nu\rho}$ satisfies, in terms of Christoffel symbols $(\Gamma^{g})_{ij}^{k}$ of the Levi-Cita connection of $g_{\mu\nu}$ (see, for example [23]),
$\Gamma_{ij}^{k}=(\Gamma^{g})_{ij}^{k}+\mathcal{A}_{i}\otimes\delta_{j}^{k}$ (38)
where $\mathcal{A}$ is a 1-form (with no restriction) and $\delta_{j}^{k}$ is Kronecker’s delta. The unique symmetric connection in (38) is the Levi-Civita one, then providing a remarkable support to standard General Relativity.
@@ 6.4.2. Palatini Approach to Finslerian Einstein Equations
$\bar{\mathscr{S}}[L,N^{\mu}_{\nu}]:=\int_{D\subset\mathbb{P}^{+}A}g^{\alpha\beta}\,\mathrm{Ric}_{\cdot\alpha\cdot\beta}\,d\Sigma^{+},$ (39)
@@ 6.4.3. Some results on the Finslerian Palatini Equations
The following sampler of results in [108] is illustrative (see more in [108, Th. D]). The first one stresses an aforementioned fundamental difference with PW.
@@
Equation (40) resembles Einstein’s, which reduced to the vanishing of the Ricci tensor just taking traces. Palatini vacuum reduces to the vanishing of the Ricci scalar (in the spirit of Rutz’s (29)) but the proof is not trivial because now a differential equation must be solved [108, Lemma 5.13]. Surprisingly, the result is obtained only for Lorentzian signature, because the proof is then based on the Riemannian maximum principle on the indicatrix [108, Th. C] (in the case that $L$ is Riemannian, an alternative proof could be given, based on the eigenvalues of the Laplacian on a round sphere, cf. [108, Th. 5.17]).
@@
In Lorentzian signature, the solutions $(L,N)$ of the Palatini equations (40), (41) such that $N$ is defined on the whole causal cone of $L$ satisfy $\mathrm{Ric}=0.$
@@
As in the case of semi-Riemannian Palatini (38), given a pseudo-Finsler metric $L$, the solutions of the affine equation (41) have a fibered structure on the set of symmetric solutions (now with fiber isomorphic to the space of anisotropic 0-homogeneous 1-forms A), [108, Th. A]. However, the uniqueness of the symmetric case is not trivial. Again, one can obtain uniqueness, by using a fiberwise argument of analiticy (notice that even non-smooth semi-Riemannian metrics are fiberwise analytic, as the metric at each point is a quadric, thus analytic). The highly original argument is valid for indefinite signature, as it will use crucially analiticity when $L$ vanishes [108, Th. B], and the solutions are assumed to be well defined therein, i.e. proper.
@@ Acknowledgments
The author acknowledges warmly D. Azagra for his advise on smoothening in the proof of Th. 3.3, F. Fernández Villaseñor (U. Castilla la Mancha) for his comments on the manuscript, especially sections 2.4.4 to 2.4.6 and 6, J. Hedicke (U. Nijmegen) for his comments on section 3.2, and M.A. Javaloyes (U. Murcia) and C. Pfeiffer (U. Bremen) for their reading of the manuscript, suggestions and references. Research partially supported by PID2024-156031NB-I00 (MCIN/ AEI /10.13039/501100011033), as well as the framework IMAG/ María de Maeztu, CEX2020-001105-MCIN/ AEI/ 10.13039/501100011033.